OpenStudy (anonymous):

how do you solve root - 9/16 with imaginary numbers?

OpenStudy (anonymous):

\[\sqrt{-9/16}\]

OpenStudy (anonymous):

and another question what is root 6 / root 3?

hero (hero):

\[\sqrt{\frac {-9}{16}} = \frac{3\sqrt{-1}}{4} = \frac {3i}{4}\]

OpenStudy (anonymous):

root(6) = root(3) * root(2) so you can factor out the root 3

OpenStudy (anonymous):

therefore root 6 / root 3 = root 2